# RIFF: Learning to Rephrase Inputs for Few-shot Fine-tuning of Language Models

### v1

@@ Abstract
Pre-trained Language Models (PLMs) can be accurately fine-tuned for downstream text processing tasks. Recently, researchers have introduced several parameter-efficient fine-tuning methods that optimize input prompts or adjust a small number of model parameters (e.g LoRA). In this study, we explore the impact of altering the input text of the original task in conjunction with parameter-efficient fine-tuning methods. To most effectively rewrite the input text, we train a few-shot paraphrase model with a Maximum-Marginal Likelihood objective. Using six few-shot text classification datasets, we show that enriching data with paraphrases at train and test time enhances the performance beyond what can be achieved with parameter-efficient fine-tuning alone.
@@ 2 Problem Formulation
$\sum_{i=1}^{N}\{\log P_{\theta_{\text{lm}}}(y_{i}|x_{i})+\\ \frac{1}{M}\sum_{j=1}^{M}\log P_{\theta_{\text{lm}}}(y_{i}|z_{i,j})\}$ (1)
@@ 2.1 LM-Friendly Paraphrase Search
Gradient Approximation: Text generation can be reformulated as an episodic reinforcement learning problem where an agent (i.e. a paraphrase generator) generates tokens (i.e. takes actions) one step at a time until reaching the end of the episode (i.e. selecting the end of sequence token). Therefore, for a given training example $(x,y)$ and its paraphrase $z$, we define the terminal reward (i.e. goodness) for $z$ as $R(z)=\log P_{\theta_{\text{lm}}}(y|z)$. When approximating the gradient vector of objective 2 concerning $\theta_{\text{par}}$, we propose two strategies. These include: (i) Maximum Marginal Likelihood (MML) and (ii) approximating the gradient vector of the paraphrase model via the Policy Gradient (PG) theorem. Notably, gradient updates using these two methods exhibit a close relationship, with the main difference lying in the posterior coefficient utilized to score each sample Guu et al. (2017). We can recast the main objective presented in equation 2 into the following equation representing the expected reward:
@@
Grammatical Integrity: We describe three distinct modeling techniques for both the MML and PG gradient estimations: On-policy learning, Off-policy learning and Proximal Policy Optimization (PPO).
@@
Our next solution for degenerate paraphrases involves imposing a penalty in the training objective if the samples drawn from the current paraphrase generator, $P_{\theta_{\text{par}}}(z|x)$, deviate from those of the pre-trained paraphrase generator. We can implement this penalty as a KL-divergence penalty between the distributions of paraphrases produced by the current model and the pre-trained one. This approach resembles the PPO learning with a KL penalty Schulman et al. (2017) and has been used in fine-tuning InstructGPT with a reward model trained over human feedback Ouyang et al. (2022). In the case of InstructGPT, researchers prevent the reward fine-tuned model from diverging from a separate language model pre-trained on supervised data Ouyang et al. (2022). To integrate this penalty we define the following new objective for $\theta_{\text{par}}$:
$J^{\text{PPO}}_{\theta_{\text{par}}}:=\log E_{z}[e^{R(z)}]-\beta E_{z}[\log s(z)]\\ s(z)=\frac{P_{\theta_{\text{par}}}(z|x)}{P_{\text{fixed}}(z|x)}$ (8)
@@
$\nabla J_{\theta_{\text{par}}}-\beta E_{z}[(\log s(z)+1)\nabla\log P_{\theta_{\text{par}}}(z|x)]\\ z\sim P_{\theta_{\text{par}}}(.|x)$ (9)
@@ 3.1 Setup
Datasets: Inspired by prior work Gao et al. (2021); Deng et al. (2022), we experiment on six classification tasks in the few-shot setting. These include sentiment classification tasks such as the binary sentiment datasets SST2 Socher et al. (2013), CR Hu and Liu (2004), and MR Pang and Lee (2005). We also experiment on the 5-label sentiment dataset SST5 Socher et al. (2013), the question type classification dataset TREC Voorhees and Tice (2000), and the topic classification dataset AGNews Zhang et al. (2015). The number of classes per dataset, as well as the used instructions are outlined in Appendix G. Instructions and class verbalizers are based on previous work Deng et al. (2022) in prompt optimization. Detailed information about the specific learning rates for each LM technique along with other hyper-parameters can be found in Appendix D.
@@ 3.2 Few-shot Paraphrase Fine-Tuing
Table 1: The average accuracy of the best performing validation checkpoint in the 128-shot SST2 classification task for both the on-policy and PPO learning techniques. Highest performance per column bolded. Last column reports the macro-average among each row. Numbers in parentheses ($\sigma$ $\mid$ $m$): $\sigma$ represents the standard deviations for the reported means across five train/validation splits; $m$ reports the average accuracy of all the validation checkpoints as we monitor robustness during the training trajectory. The suffix ‘-$Z$’ denotes models trained with reward normalization.
Learning On-Policy PPO AVG
@@
We assess the impact of reward normalization in the context of on-policy, off-policy, and PPO learning, considering both PG and MML gradient estimations. Table 1 lists the best performance out of all the checkpoints evaluated on the validation splits, which is further averaged over five validation splits. With both PG and MML gradient estimations, reward normalization is boosting the performance across the three text decoding techniques for both on-policy and PPO learning techniques (see ‘AVG’ column in Table 1). Conversely, reward normalization is not improving performance with off-policy learning (follow discussion in Appendix B and see Table 4)
Table 1 verifies that MML gradient estimation outperforms PG gradient estimation on average across three decoding techniques for both on-policy and PPO learning techniques. The highest accuracy is achieved by ‘PG-Z’ with on-policy learning and top-p decoding, however it is not robust during the entire training trajectory. Figure 1 shows that PG gradient estimation is not robust throughout the training trajectory, which causes the paraphrase generator to produce nonsensical paraphrases. This results in downstream classification performance on par with random guessing. In contrast, off-policy and PPO learning circumvent this divergence. MML gradient estimation maintains robustness throughout the training phase. In Table 1, we also report the average accuracy of all the checkpoints as we are training the models (numbers in parentheses). The learning technique ‘MML-Z’ is more robust during the training trajectory compared to ‘PG-Z’.
Upon investigating various elements of our learning objectives for fine-tuning the paraphrase generator, the combination that delivers the best performance across the validation splits, which is also robust during the entire training trajectory, includes: MML gradient approximation, PPO learning, mixed decoding for sample generation, and finally applying reward normalization. We name this combined approach our proposed RIFF algorithm.
@@ 3.3 Paraphrase Quality Analysis
Semantic Similarity ($SE$): To assess how semantically similar paraphrases are to the original input text, we compute the BERTScore’s F1 metric Zhang* et al. (2020) between each paraphrase and the original input text using the BERT-Large model. A higher $SE$ score signifies more semantic similarity with the original text.
@@
Table 2: Average metrics on the test sets to evaluate the quality of paraphrases across three classification tasks: SST2, SST5, and AGNews datasets. The metrics are further averaged across five training folds for the FinPara method. OrigIn: Represents the original task inputs. PrePara: Corresponds to task inputs obtained from the pre-trained paraphraser. FinPara: Indicates task inputs from the finetuned paraphraser in the $128$-shot setting. We scale scores into the range of [0-100], except for GRGR. Better performance per column bolded
Input Type GRGR LDLD PLDPLD SESE FCFC
@@
We present the metrics for the datasets in Table 2. Compared to a model that was not fine-tuned for this task (PrePara), our RIFF algorithm has reduced the diversity among the generated paraphrases and their lexical variation compared to the original input text. This outcome aligns with our search-learn objective that prioritizes high-scoring paraphrases over others. RIFF has contributed to higher semantic similarity compared to the original input. Interestingly, the perplexity of the generated paraphrases after fine-tuning with our objective is still low, demonstrating the grammatical accuracy of these paraphrases. The example paraphrases shown in Appendix C illustrate that RIFF reduces hallucination in the generated paraphrases, which may contribute to the lower $LD$ score but higher $SE$ score with respect to the original input text. A higher $FC$ score as shown in Table 2 verifies that the RIFF objective has reduced hallucination in the generated paraphrases.
@@ 3.4 Paraphrases for Few-shot LM Tuning
Table 3: Average accuracy on the standard evaluation sets for the 16-shot text classification. Numbers in parentheses are standard deviations across the five train/validation folds. The last column is the micro averaged performance across the datasets. Highest performance per dataset bolded, second highest underlined. $\dagger$: the average 16-shot AllTune results with automatically searched templates Gao et al. (2021). $\star$: reported zero-shot results for GPT3 with in-context learning Gao et al. (2021).
@@
Table 3 illustrates the average accuracy on standard test sets across six text classification datasets. The reported scores correspond to four distinct LM tuning techniques: GS, SpTune, AllTune, and LoRA. Results for ClsTune, HTune, and InTune are presented in Table 11 of Appendix E.
@@ 3.5 Paraphrase Robustness Analysis
We generate $M=8$ paraphrases for each test input text and investigate the average performance of the fine-tuned language models on each of these paraphrases compared to the original input text. Figure 2 illustrates the average test accuracy over six classification datasets for AllTune, LoRA, and SPTune, which are popular LM tuning techniques in NLP. The original input texts are denoted with inputs at rank 0, whereas rank_ii ($i\in\{1,2,3,4,5,6,7,8\}$) are the paraphrases returned by the diverse beam search. We observe that the LM tuning techniques are not robust to paraphrases. Our paraphrase augmentation during training increases paraphrase robustness. SPTune observes more significant improvement in robustness from paraphrase augmentation during training.
@@ 4 Related Works
Prompt Optimization & Efficient Tuning: Recent research proposes various techniques for prompt optimization and efficient tuning. We have used successful techniques from each of these areas. Appendix F provides our detailed description of these recent techniques. All of the recent techniques for prompt optimization and efficient tuning use the original input context provided within the dataset.
Paraphrase Generation: Recent techniques encompass various approaches, including the use of copy mechanisms, Variational Autoencoders, Generative Adversarial Networks, and Reinforcement Learning techniques to generate diverse paraphrases Zhou and Bhat (2021). While previous studies have applied RL techniques for paraphrase generation, we propose the use of MML gradients instead of policy gradients to fine-tune our paraphrase model by the reward of a secondary classification task (see Appendix F for more discussion).
@@ 5 Conclusion
We investigated the impact of incorporating input paraphrases while fine-tuning PLMs with recent efficient tuning techniques. We also provided extensive experiments for reducing noise in a distantly supervised paraphrase generator.
Limitations
We hypothesize that our proposed objective can be used to generate paraphrases that minimize the downstream language model’s performance on unseen data by maximizing the objective $\log(1-P(y|x))$. Exploring the connection to Adversarial Generative Networks and textual attacks on language models through paraphrase generation presents a promising avenue for future work.
@@ Appendix A Baseline LM Tuning Techniques
Classifier-Finetuning (ClsTune): In ClsTune, we first create a feature representation $h(x)$ for the input text $x$ using average pooling of the final hidden vectors in the last layer of the language model. Here, we assume that the language model (feature extractor) remains fixed, and we then construct a two-layer feedforward network with the gelugelu activation function Hendrycks and Gimpel (2016) as a classification module on top of the language model.
@@ Appendix D Further Training Details
Paraphrase sample size MM=8
@@
Appendix E Paraphrases for Few-shot LM Tuning (Further Results)
Due to space limitations, we present the results for ClsTune, HTune, and InTune in Table 11.
Table 11: Average accuracy on the standard evaluation sets for the 16-shot text classification using the ClsTune, HTune, and InTune fine-tuning techniques. Numbers in parentheses are standard deviations across the five train/validation folds. The last column is the micro averaged performance across the datasets.
Tuning Method SST2 SST5 CR MR TREC AGN AVG
Appendix F Extended Related Works
@@
Appendix G Task Instructions & Input Format

### v2

@@ Abstract
Pre-trained Language Models (PLMs) can be accurately fine-tuned for downstream text processing tasks. Recently, researchers have introduced several parameter-efficient fine-tuning methods that optimize input prompts or adjust a small number of model parameters (e.g LoRA). In this study, we explore the impact of altering the input text of the original task in conjunction with parameter-efficient fine-tuning methods. To most effectively rewrite the input text, we train a few-shot paraphrase model with a Marginal Maximum Likelihood objective. Using six few-shot text classification datasets, we show that enriching data with paraphrases at train and test time enhances the performance beyond what can be achieved with parameter-efficient fine-tuning alone. The code used for our experiments can be found at https://github.com/SaeedNajafi/RIFF.
@@ 2 Problem Formulation
$\sum_{i=1}^{N}\Bigl{\{}\log P_{\theta_{\text{lm}}}(y_{i}|x_{i})+\\ \frac{1}{M}\sum_{j=1}^{M}\log P_{\theta_{\text{lm}}}(y_{i}|z_{i,j})\Bigr{\}}$ (1)
@@ 2.1 LM-Friendly Paraphrase Search
Gradient Approximation: Text generation can be reformulated as an episodic reinforcement learning problem where an agent (i.e. a paraphrase generator) generates tokens (i.e. takes actions) one step at a time until reaching the end of the episode (i.e. selecting the end of sequence token). Therefore, for a given training example $(x,y)$ and its paraphrase $z$, we define the terminal reward (i.e. goodness) for $z$ as $R(z)=\log P_{\theta_{\text{lm}}}(y|z)$. When approximating the gradient vector of objective 2 concerning $\theta_{\text{par}}$, we propose two strategies. These include: (i) Marginal Maximum Likelihood (MML) and (ii) approximating the gradient vector of the paraphrase model via the Policy Gradient (PG) theorem. Notably, gradient updates using these two methods exhibit a close relationship, with the main difference lying in the posterior coefficient utilized to score each sample Guu et al. (2017). We can recast the main objective presented in equation 2 into the following equation representing the expected reward:
@@
Grammatical Integrity: We describe three distinct modeling techniques for both the MML and PG gradient estimations: On-policy learning, Off-policy learning and KL-penalized On-policy (KLOn) learning.
@@
Our next solution for degenerate paraphrases involves imposing a penalty in the training objective if the samples drawn from the current paraphrase generator, $P_{\theta_{\text{par}}}(z|x)$, deviate from those of the pre-trained paraphrase generator. We can implement this penalty as a KL-divergence penalty between the distributions of paraphrases produced by the current model and the pre-trained one. To integrate this penalty we define the following new objective for $\theta_{\text{par}}$:
$J^{\text{KLOn}}_{\theta_{\text{par}}}:=\log E_{z}[e^{R(z)}]-\beta E_{z}[\log s(z)]$ (8)
@@
$\nabla J_{\theta_{\text{par}}}-\beta E_{z}[(\log s(z)+1)\nabla\log P_{\theta_{\text{par}}}(z|x)],\\ z\sim P_{\theta_{\text{par}}}(.|x)$ (9)
@@ 3.1 Setup
Datasets: Inspired by prior work Gao et al. (2021); Deng et al. (2022), we experiment on six classification tasks in the few-shot setting. These include sentiment classification tasks such as the binary sentiment datasets SST2 Socher et al. (2013), CR Hu and Liu (2004), and MR Pang and Lee (2005). We also experiment on the 5-label sentiment dataset SST5 Socher et al. (2013), the question type classification dataset TREC Voorhees and Tice (2000), and the topic classification dataset AGNews Zhang et al. (2015). The number of classes per dataset, as well as the used instructions are outlined in Appendix F. Instructions and class verbalizers are based on previous work Deng et al. (2022) in prompt optimization. Detailed information about the specific learning rates for each LM technique along with other hyper-parameters can be found in Appendix D.
@@ 3.2 Few-shot Paraphrase Fine-Tuing
Table 1: The average accuracy of the best performing validation checkpoint in the 128-shot SST2 classification task for both the on-policy and KLOn learning techniques. Highest performance per column bolded. Last column reports the macro-average among each row. Numbers in parentheses ($\sigma$mid $m$): $\sigma$ represents the standard deviations for the reported means across five train/validation splits; $m$ reports the average accuracy of all the validation checkpoints as we monitor robustness during the training trajectory. The suffix ‘-$Z$’ denotes models trained with reward normalization.
Learning On-Policy KLOn AVG
@@
We assess the impact of reward normalization in the context of on-policy, off-policy, and KL-penalized on-policy (KLOn) learning, considering both PG and MML gradient estimations. Table 1 lists the best performance out of all the checkpoints evaluated on the validation splits, which is further averaged over five validation splits. With both PG and MML gradient estimations, reward normalization is boosting the performance across the three text decoding techniques for both on-policy and KLOn learning techniques (see ‘AVG’ column in Table 1). Conversely, reward normalization is not improving performance with off-policy learning (follow discussion in Appendix B and see Table 4)
Table 1 verifies that MML gradient estimation outperforms PG gradient estimation on average across three decoding techniques for both on-policy and KLOn learning techniques. The highest accuracy is achieved by ‘PG-Z’ with on-policy learning and top-p decoding, however it is not robust during the entire training trajectory. Figure 1 shows that PG gradient estimation is not robust throughout the training trajectory, which causes the paraphrase generator to produce nonsensical paraphrases. This results in downstream classification performance on par with random guessing. In contrast, off-policy and KLOn learning circumvent this divergence. MML gradient estimation maintains robustness throughout the training phase. In Table 1, we also report the average accuracy of all the checkpoints as we are training the models (numbers in parentheses). The learning technique ‘MML-Z’ is more robust during the training trajectory compared to ‘PG-Z’.
Upon investigating various elements of our learning objectives for fine-tuning the paraphrase generator, the combination that delivers the best performance across the validation splits, which is also robust during the entire training trajectory, includes: MML gradient approximation, KLOn learning, mixed decoding for sample generation, and finally applying reward normalization. We name this combined approach our proposed RIFF algorithm.
@@ 3.3 Paraphrase Quality Analysis
Semantic Similarity ($SS$): To assess how semantically similar paraphrases are to the original input text, we compute the BERTScore’s F1 metric Zhang* et al. (2020) between each paraphrase and the original input text using the BERT-Large model. A higher $SS$ score signifies more semantic similarity with the original text.
@@
Table 2: Average metrics on the test sets to evaluate the quality of paraphrases across three classification tasks: SST2, SST5, and AGNews datasets. The metrics are further averaged across five training folds for the FinPara method. OrigIn: Represents the original task inputs. PrePara: Corresponds to task inputs obtained from the pre-trained paraphraser. FinPara: Indicates task inputs from the finetuned paraphraser in the $128$-shot setting. We scale scores into the range of [0-100], except for $GR$. Better performance per column bolded.
Input Type $GR$ $LD$ $PLD$ $SS$ $FC$
@@
We present the metrics for the datasets in Table 2. Compared to a model that was not fine-tuned for this task (PrePara), our RIFF algorithm has reduced the diversity among the generated paraphrases and their lexical variation compared to the original input text. This outcome aligns with our search-learn objective that prioritizes high-scoring paraphrases over others. RIFF has contributed to higher semantic similarity compared to the original input. Interestingly, the perplexity of the generated paraphrases after fine-tuning with our objective is still low, demonstrating the grammatical accuracy of these paraphrases. The example paraphrases shown in Appendix C illustrate that RIFF reduces hallucination in the generated paraphrases, which may contribute to the lower $LD$ score but higher $SS$ score with respect to the original input text. A higher $FC$ score as shown in Table 2 verifies that the RIFF objective has reduced hallucination in the generated paraphrases.
@@ 3.4 Paraphrases for Few-shot LM Tuning
Table 3: Average accuracy on the standard evaluation sets for the 16-shot text classification. Numbers in parentheses are standard deviations across the five train/validation folds. The last column is the micro averaged performance across the datasets. Highest performance per dataset bolded, second highest underlined. $\dagger$: the average 16-shot AllTune results with automatically searched templates Gao et al. (2021). $\star$: reported results for RoBERTa-large using in-context learning Gao et al. (2021).
@@
Table 3 illustrates the average accuracy on standard test sets across six text classification datasets. The reported scores correspond to seven distinct LM tuning techniques: ClsTune, GS, SpTune, HTune, InTune, AllTune, and LoRA.
@@ 3.5 Paraphrase Robustness Analysis
We generate $M=8$ paraphrases for each test input text and investigate the average performance of the fine-tuned language models on each of these paraphrases compared to the original input text. Figure 2 illustrates the average test accuracy over six classification datasets for AllTune, LoRA, and SPTune, which are popular LM tuning techniques in NLP. The original input texts are denoted with inputs at rank 0, whereas rank_$i$ ($i\in\{1,2,3,4,5,6,7,8\}$) are the paraphrases returned by the diverse beam search. We observe that the LM tuning techniques are not robust to paraphrases. Our paraphrase augmentation during training increases paraphrase robustness. SPTune observes more significant improvement in robustness from paraphrase augmentation during training.
3.6 Limitations Discussion
In terms of training overhead for our method, once the paraphrase model is fine-tuned, augmenting the mini-batches with paraphrases has minimal impact on training time in the downstream classification task. This is because we only need to perform inference with the paraphrase model once for the training examples to generate their paraphrases before the first epoch. The generated paraphrases are then cached for subsequent training epochs while fine-tuning the downstream language model.
In contrast, fine-tuning the paraphrase model using the RIFF objective requires generating new paraphrases each epoch, which prevents us from caching paraphrases across epochs. This fine-tuning step is about eight times slower than maximum-likelihood training, which benefits from having ground-truth paraphrases available. Generating new samples and scoring them is a standard procedure in Reinforcement Learning.
@@ 4 Related Works
Prompt Optimization & Efficient Tuning: Recent research proposes various techniques for prompt optimization and efficient tuning. We have used successful techniques from each of these areas. Appendix E provides our detailed description of these recent techniques. All of the recent techniques for prompt optimization and efficient tuning use the original input context provided within the dataset.
Paraphrase Generation: Recent techniques encompass various approaches, including the use of copy mechanisms, Variational Autoencoders, Generative Adversarial Networks, and Reinforcement Learning techniques to generate diverse paraphrases Zhou and Bhat (2021). While previous studies have applied RL techniques for paraphrase generation, we propose the use of MML gradients instead of policy gradients to fine-tune our paraphrase model by the reward of a secondary classification task (see Appendix E for more discussion).
@@ 5 Conclusion
We investigated the impact of incorporating input paraphrases while fine-tuning PLMs with recent efficient tuning techniques. Our results indicate that specific techniques, such as continuous and discrete prompt optimization methods like AutoPrompt or Soft-Prompt Tuning, benefit significantly from the inclusion of paraphrases. We also conducted extensive experiments to reduce noise in a distantly supervised paraphrase generator. Our ablation studies on fine-tuning the paraphrase generator demonstrate that policy gradient objectives lack robustness during training, while maximum marginal likelihood training remains a robust technique.
@@
Acknowledgements
We gratefully acknowledge the Vector Institute, and the Digital Research Alliance of Canada (alliancecan.ca) for providing access to their GPU cluster, which streamlined our experiments. This research was funded by NSERC, Alberta Machine Intelligence Institute (Amii), and CIFAR. Alona Fyshe is a Canada CIFAR AI Chair.
@@ Appendix A Baseline LM Tuning Techniques
Classifier-Finetuning (ClsTune): In ClsTune, we first create a feature representation $h(x)$ for the input text $x$ using average pooling of the final hidden vectors in the last layer of the language model. Here, we assume that the language model (feature extractor) remains fixed, and we then construct a two-layer feedforward network with the $gelu$ activation function Hendrycks and Gimpel (2016) as a classification module on top of the language model.
@@ Appendix D Further Training Details
Paraphrase sample size $M$=8
@@
Appendix E Extended Related Works
@@
Appendix F Task Instructions & Input Format